### v2

# Breaking the Assistant Mold: Modeling Behavioral Variation in LLM Based Procedural Character Generation

Abstract
Procedural content generation has enabled vast virtual worlds through levels, maps, and quests, but large-scale character generation remains underexplored. We identify two alignment-induced biases in existing methods: a positive moral bias, where characters uniformly adopt agreeable stances (e.g., always saying lying is bad), and a helpful assistant bias, where characters invariably answer questions directly (e.g., never refusing or deflecting). While such tendencies suit instruction-following systems, they suppress dramatic tension and yield predictable characters, stemming from maximum likelihood training and assistant fine-tuning. To address this, we introduce PersonaWeaver, a framework that disentangles world-building (roles, demographics) from behavioral-building (moral stances, interactional styles), yielding characters with more diverse reactions and moral stances, as well as second-order diversity in stylistic markers like length, tone, and punctuation. Code: https://github.com/mqraitem/Persona-Weaver
Breaking the Assistant Mold: Modeling Behavioral Variation in LLM Based Procedural Character Generation
Maan Qraitem, Kate Saenko, Bryan A. Plummer Boston University {mqraitem, saenko, bplum}@bu.edu
1 Introduction
Figure 1: Illustration of biases in prior work versus our method Prior work populations Jin et al. (2024) in (a)-Top collapse into uniform agreement with moral statements from the Social Chemistry dataset Forbes et al. (2020), reflecting a moral bias, while in (a)-Bottom they invariably answer conversational questions from ConvAI2 Dinan et al. (2019), reflecting a reaction bias. In contrast, our generated populations (b)-Top and (b)-Bottom exhibit a broader range of stances and responses.
Procedural character generation (PCG) is a recent yet underexplored task that aims to automatically create diverse and believable agents for virtual and narrative environments. Enabled by advances in large language models (LLMs), which unlock new possibilities for character role play Wang et al. (2023); Zhou et al. (2023); Shao et al. (2023), LLMs are an attractive foundation for PCG. Existing approaches—either directly prompting LLMs to produce populations Jin et al. (2024) or adapting scraped persona banks to specific settings Ge et al. (2024)—inherit alignment-induced biases: a positive moral bias, where characters uniformly adopt agreeable stances (Fig. 1 Top), and a helpful assistant bias, where they invariably answer questions (Fig. 1 Bottom). While such tendencies are desirable in instruction-following systems, they suppress dramatic tension and lead to repetitive archetypes. This reflects a broader limitation of maximum-likelihood training and assistant fine-tuning, which bias models toward safe continuations and homogenized voices, as also observed in other simulation tasks Kotek et al. (2023); Cheng et al. (2023); Wang et al. (2025).
Figure 2: Overview of PersonaWeaver, which disentangles world-building (a) from behavioral modeling (b) to give explicit control over behavioral variation (Moral Attitudes in this instance). A final Sample and Mix step (c) then combines these components into character profiles and ensures variation. Refer to Section 2.3 for further discussion.
To address these issues, we introduce PersonaWeaver, which tackles the collapse of behavior by disentangling world-building from behavioral-building (Fig. 2). This separation allows us to explicitly model variation in behavior, providing finer control to probe whether LLMs can realize diverse dispositions. Unlike world-building, which is setting-specific (e.g., a corn farmer does not live in New York City), behavioral traits such as moral stances and interactional styles are universal, making them amenable to external banks that allow systematic specification. For example, we use Moral Foundations Theory Graham et al. (2013) to define stances across axes such as care, fairness, and authority, and an interactional bank to capture conversational tendencies such as cooperation, disagreement, or evasion. By sampling and recombining these dimensions, PersonaWeaver creates character profiles that exhibit varied reactions and moral stances (Fig. 2). In sum, our results demonstrate that PersonaWeaver elicits substantially more diverse responses than prior work when behavior is explicitly modeled (Fig. 1 b).
We demonstrate this through a large-scale experimental procedure, instantiating character profiles across fictional and non-fictional settings and testing them on three frontier models: GPT-4o Achiam et al. (2023), LLaMA 4 AI (2024), and Qwen 3 Yang et al. (2025). Characters are then probed on two tasks. In a moral belief experiment adapted from Social Chemistry Forbes et al. (2020), they are presented with normative statements (e.g., , Fig. 1 a) and asked to register agreement or disagreement; PersonaWeaver yields a broader, more balanced distribution of stances than prior work, avoiding collapse to uniform agreement. In an interaction experiment using conversational prompts from ConvAI2 Dinan et al. (2019), characters produce open-ended responses; here, PersonaWeaver exhibits more varied behaviors—refusal, deflection, disagreement—while also inducing second-order stylistic diversity in features such as length, punctuation, and emotional tone, resulting in richer and less homogenized profiles.
Our contributions are twofold:
  • We identify two alignment-induced biases in existing methods for PCG: a positive moral bias and a helpful assistant bias.
  • We propose PersonaWeaver which disentangles world-building from behavioral-building and mitigates the aforementioned biases while generating diverse behaviors.
2 A Tale of Two Biases in PCG
Given a setting $T$ (e.g., rural town), we define Procedural Character Generation as sampling a population of textual descriptions $P=\{d_{1},\dots,d_{n}\}$, where each $d_{i}$ describes a character in $T$. These descriptions condition an LLM $f_{\theta}$, which simulates interactive agents $c_{i}=f_{\theta}(d_{i},T)$ that act in context as characters.
We are concerned with identifying and mitigating two types of behavioral biases that arise in existing approaches to LLM-based procedural character generation Jin et al. (2024); Ge et al. (2024): (1) positive moral bias, where characters overwhelmingly agree on normative statements, and (2) helpful assistant bias, where characters invariably answer questions directly. To this end, we define our study preliminaries in Sec. 2.1, empirically demonstrate the two biases in Sec. 2.2, and introduce PersonaWeaver in Sec. 2.3 and show how it mitigates them.
2.1 Preliminaries
Prior Work. WorldWeaver Jin et al. (2024) generates characters through direct prompting of the LLM (e.g., Generate $N$ different Character profiles for Setting $T$). PersonaHub Ge et al. (2024) samples personas from an internet-scale bank of scraped profiles and adapts them to the target setting using an LLM.
Simulations. Prior work Jin et al. (2024); Ge et al. (2024) has evaluated character generation within a single setting, limiting insight into how methods generalize across diverse contexts. Thus, we construct a broader experimental environment by instantiating populations across 10 distinct settings (five fictional, five real-life), sampled from popular movies and television to ensure cultural and geographic diversity. For each setting, we generate 100 characters, yielding 1,000 characters per method. Refer to Appendix E for further details.
Figure 3: Comparison of Moral Positions between prior work WorldWeaver Jin et al. (2024), PersonaHub Ge et al. (2024) and our method PersonaHub. Refer to Sec 2.2 and 2.3.1 for discussion.
Figure 4: Comparison of Reactions to Questions between prior work WorldWeaver Jin et al. (2024), PersonaHub Ge et al. (2024) and our method PersonaHub. Refer to Sec 2.2 and 2.3.1 for discussion.
Quantifying Biases. We examine moral bias, where characters uniformly adopt agreeable stances, and reaction bias, where they invariably answer questions directly. To probe moral bias, Per Fig. 1 Top, we present characters with moral norm statements from Social Chemistry Forbes et al. (2020) and measure the distribution of multiple-choice responses. To probe reaction bias, Per Fig. 1 Bottom, we ask characters conversational questions drawn from ConvAI2 Dinan et al. (2019) ; their open-ended replies are classified by an auxiliary LLM into three categories (refusal, deflection, compliance), yielding distributions over behaviors. Refer to Appendix B for further details.
Models. Our results in the paper are based on GPT-4o Achiam et al. (2023), a frontier LLM. To ensure generality, we replicate all experiments on LLaMA 4 AI (2024) and Qwen 3 Yang et al. (2025); in Appendix F.
(a) Filler Words
(b) Punctuations
(c) Answer Length
(d) Sentiment
Figure 5: Comparison of Stylistic Patterns in the generated answers of prior work (WorldWeaver Jin et al. (2024) and PersonaHub Ge et al. (2024)) and our work PersonaWeaver across four Stylistic categories (Filler words, Punctuations, Answer Length, and Sentiment). Refer to Section 2.3.1 for further discussion.
2.2 Demonstrating Biases Empirically
Moral Bias. Fig. 3 shows that prior methods such as WorldWeaver and PersonaHub overwhelmingly default to agreement when judging normative statements. Characters consistently adopt positive, prosocial stances, yielding homogenized populations rather than a spectrum of moral positions.
Reaction Bias. As shown in Fig. 4, WorldWeaver and PersonaHub characters almost invariably comply with conversational prompts, directly answering questions rather than refusing, deflecting, or resisting. This collapse into uniform compliance prevents the range of evasive or resistant reactions characteristic of natural dialogue.
Discussion. These results show how alignment-driven objectives—maximum likelihood estimation, which favors high-probability continuations, and assistant fine-tuning Ouyang et al. (2022), which rewards helpfulness—collapse behavioral diversity. Characters thus inherit a moral bias toward universal agreement and a reaction bias toward always answering, reinforcing predictable, assistant-like voices. In Sec. 2.3, we introduce PersonaWeaver, designed to probe and mitigate these biases.
2.3 PersonaWeaver: Disentangling World from Behavior Building
PersonaWeaver addresses the collapse of behavior by explicitly disentangling world-building from behavioral modeling. The role of this disentanglement is to isolate behavioral traits from world attributes, giving us fine-grained control over how variation is introduced. To that end, we introduce three components below: a world-building module, a behavioral module, and a Sample and Mix module that combines attributes from both.
World-Building Module. Per Fig. 2, to prevent behavioral leakage into world attributes, we begin by asking the model to propose axes of variation such as occupation, affiliation, or expertise, with explicit instructions to exclude behavioral traits. Each axis is then expanded into options tailored to the specific setting (e.g., , “farmer” in a rural town), producing a bank of non-behavioral attributes from which character profiles can be sampled.
Behavioral Module. Unlike world-building, which is setting-dependent, behavioral traits are universal, so we define them through external banks that give us systematic control. For moral traits, we draw inspiration from Moral Foundations Theory Graham et al. (2013), defining eight distinct moral stances that are sampled and attached to character profiles, producing systematic variation in how characters judge social norms (full list in the Appendix C). For interactional traits, we define eight categories spanning a spectrum of conversational behaviors: refusals, deflections, hesitations, direct compliance, volunteering additional details, playful or subversive replies, hostile retorts, and meta-level commentary. Together, these banks allow us to break the “always answer” defaults that dominate prior work. Refer to Appendix D for further details.
Sample and Mix. The module recombines axes from both the world-building and behavioral modules in a Sample and Mix step (Fig. 2). By randomly mixing attributes across dimensions, we force unlikely combinations that prevent collapse into narrow archetypes and systematically test whether LLMs can sustain diverse behaviors once specified. Finally, we prompt the LLM to flag implausible pairings (e.g., , “a medieval knight software engineer”) and minimally revise them for coherence without reintroducing uniformity..
2.3.1 PersonaWeaver Mitigates Biases
Broader Moral Coverage. Per Fig. 3, PersonaWeaver yields a more balanced distribution of moral stances than prior work. Instead of collapsing to near-universal agreement, populations generated with PersonaWeaver spread across the full range of options, capturing disagreement and ambivalence alongside prosocial agreement.
Broader Reaction Coverage. Per Fig. 4, PersonaWeaver significantly increases the variety of conversational behaviors. Whereas prior work overwhelmingly defaults to compliance, characters generated with PersonaWeaver exhibit refusals and deflections. This shows that our behavioral bank successfully counteracts the helpful-assistant bias.
Stylistic Variation. Finally, we analyze second-order stylistic effects (Fig. 5a-c). Characters generated with PersonaWeaver not only differ in moral and interactional stances, but also in expressive style: contain different rates of filler words (a) employ more diverse punctuation (b) and their responses vary more in length (c). These emergent differences suggest that once primary behavioral variation is enforced, LLMs naturally extend this into stylistic dimensions, producing populations that are richer and less homogenized overall.
Sentiment Diversity. Per Fig. 5 (d), PersonaWeaver broadens the sentiment distribution of character responses. Whereas prior work clusters around uniformly positive sentiment, PersonaWeaver yields a richer emotional spread—including neutral and negative tones—that better reflects the real-world.
3 Conclusion
We showed that LLM-based character generation exhibit moral and reaction biases. Thus, we introduced PersonaWeaver which mitigates said biases by disentangling world-building from behavioral modeling and using external banks to impose explicit behavioral variation. Experiments reveal more diverse reactions, moral stances, and second-order stylistic richness, demonstrating a more expressive range of LLM-generated characters.
4 Limitations
Out evaluation is limited in two major ways: 1) it only examines two behavioral dimensions: moral stances and characters’ interactions. It overlooks dimensions like emotional regulation. Therefore, future work can benefit from expanding our evaluation setup 2) While our coarse grained moral stance evaluation enables a streamlined study, it overlooks more nuanced moral reasoning that can’t be simply coarsely categorized.
Potential Risks Efforts to systemically expand behavioral diversity in character generation can push models to generate characters that replicate offensive or unsafe behaviors. Therefore, careful curation of behavioral banks is essential to ensure that increasing diversity in character generation serves creative and research goals without amplifying harm.
Appendix A Related Works
Procedural Character Generation. Research on procedural character generation remains limited, with most prior work focusing on producing characters within a single, predefined environment (e.g., , one game world Jin et al. (2024); Ge et al. (2024)). WorldWeaver Jin et al. (2024) prompts an LLM to directly generate $N$ characters for a given setting, while PersonaHub Ge et al. (2024) samples personas from large-scale scraped profile datasets and adapts them to the target context using an LLM. In this work, we show that these approaches struggle with two key issues—limited moral diversity and restricted interactional range—and introduce a method designed to explicitly mitigate these biases.
Biases in Simulating Personas. Prior studies have documented a broader lack of behavioral and identity diversity in simulated personas. For instance, Marked Personas Cheng et al. (2023) show that language models tend to reproduce social stereotypes, while others highlight gender and identity flattening Kotek et al. (2023); Wang et al. (2025) and reduced narrative variety in multimodal storytelling Lee and Jeon (2024). Such homogenization arises from maximum-likelihood training objectives that favor high-probability continuations, as well as alignment fine-tuning that rewards politeness, safety, and helpfulness—thereby suppressing the variability essential to believable social simulation. In our work, we extend this observation to procedural character generation, showing that similar alignment-induced biases constrain moral and interactional diversity in LLM-generated populations.
Character Simulation and Role-Playing. A growing line of research explores how LLMs can function as conversational agents with consistent personality, memory, and long-term coherence. Generative Agents Park et al. (2023) simulate memory-driven individuals inhabiting a shared environment, though their backgrounds are largely hand-crafted. In the role-playing domain, works such as RoleLLM Wang et al. (2023), CharacterGLM Zhou et al. (2023), and Character-LLM Shao et al. (2023) focus on eliciting and sustaining role-specific behaviors through persona-conditioned dialogue systems. These efforts primarily target the believability and consistency of agent role play. In contrast, our work examines whether off-the-shelf LLMs can leverage their broad world knowledge to simulate behaviorally diverse populations—spanning distinct moral dispositions and interactional tendencies—rather than a few coherent yet homogeneous individuals.
Source Questions / Statements
Social Chemistry Forbes et al. (2020) Parents are expected to make sure their kids eat healthy food.
It’s rude to fail to recognize other’s contributions.
You shouldn’t be noisy around others so you don’t disturb them.
It’s good to have faith in your skills as a performer.
It’s kind to let other trust their judgment.
It is bad to make lies.
It’s important to be a team player when playing team sports.
You should not steal things from other people.
You shouldn’t disrespect other people.
It’s bad to make assumptions about people based on their looks.
ConvAI2 Dinan et al. (2019) Sentiment (s1–s5):
How are you feeling?
What’s on your mind?
What else is worrying you?
What are you excited about?
What are you anxious about?
General (q1–q5):
What do you do for fun?
What’s your dream job?
What is your favorite thing to do with your family?
What kind of music do you like to play?
What’s your favorite animal?
Table 1: Moral statements and conversational questions used in our evaluation. Moral norms are drawn from Social Chemistry Forbes et al. (2020), while conversational prompts are extracted from ConvAI2 Dinan et al. (2019). Refer to Appendix B for further discussion.
ID Moral Position
M1 Protect others from harm, show compassion. Be fair and honest, avoid cheating. Be loyal to friends and allies. Question authority and resist control. Ignore purity or sacred rules. Defend personal freedom and resist oppression.
M2 Help others only when it’s easy or benefits you. Be fair when it serves your interests, otherwise bend the rules. Support your group if it gives you an advantage, but walk away if not. Obey authority when it benefits you, ignore it when it doesn’t. Dismiss purity or sacred values as irrelevant. Accept limits on freedom if they bring you personal gain.
M3 Show kindness but prioritize your own circle. Be fair within your group, less so with outsiders. Put family, community, or team first. Respect tradition and leaders. Value purity and avoid what feels degrading. Do not challenge authority for the sake of freedom.
M4 Ignore others’ suffering, focus only on your own comfort. Cheat and deceive if it benefits you. Abandon groups and betray others when useful. Reject rules, defy authority, and act for yourself. Mock sacred things and indulge in what others call impure. Dominate others and accept domination if it helps you survive.
M5 Show compassion but avoid being naive. Be fair, but allow exceptions. Be cooperative when it benefits the team. Respect some rules but question others. Avoid disgusting or degrading acts. Seek personal freedom but balance with order.
M6 Prioritize kindness and protect the vulnerable. Defend justice and equal treatment for all. Support groups but not blindly. Challenge unfair authority. Treat purity concerns as symbolic, not binding. Strongly defend freedom and resist control.
M7 Be polite and respectful to others. Be truthful and fair. Act as a dependable teammate. Obey rules and respect tradition. Honor sacred or cultural norms. Accept limits on freedom for social order.
M8 Show occasional kindness but focus on yourself. Bend rules when you can get away with it. Switch loyalty depending on advantage. Undermine authority if it benefits you. Ignore purity rules unless convenient. Value freedom only when it serves you.
Table 2: The eight moral positions used in our evaluation, inspired by Moral Foundations Theory Graham et al. (2013). Each position encodes a distinct stance toward compassion, fairness, loyalty, authority, purity, and liberty. Refer to Appendix C for further discussion.
ID Reaction Category
0 Refusal — Refuses to answer the question.
1 Deflection — Redirects or dismisses the question.
2 Hesitation — Hesitates about whether to answer.
3 Compliance — Answers the question directly.
4 Volunteering — Answers fully and adds extra details, even unasked ones.
5 Playful/Subversive — Replies in a teasing, sarcastic, or ironic way.
6 Hostile — Responds with aggression, sarcasm, or dismissal of the asker.
7 Meta — Comments on the act of being questioned itself instead of answering.
Table 3: Interactional reaction categories used in the behavioral module of PersonaWeaver. These categories expand the expressive space beyond the default “always answer” bias.
Realistic Settings Prompt
Friday Night Lights A rural Texas town with a high school football stadium, modest houses, and wide flat plains.
Seinfeld A Manhattan neighborhood block with apartment buildings, cafes, and subway entrances on busy city streets.
Fargo A Midwestern town in Minnesota with snow-covered roads, low-rise shops, and roadside diners.
The Wire A Baltimore neighborhood with brick townhouses, narrow alleys, and corner stores.
Lady Bird Sacramento with a Catholic school campus, residential streets, and modest houses.
Fantastical Settings Prompt
Wizard of Oz A fantastical city with tall green towers and glittering walls, inhabited by magical beings and travelers from distant lands.
Frozen A Nordic-inspired kingdom with a fjord-side castle, alpine peaks, and snow-covered villages, inhabited by royal families and townspeople.
Game of Thrones A medieval coastal city with high stone walls, winding streets, and a fortress overlooking the harbor, inhabited by nobles, soldiers, and commoners.
Avatar An alien moon with towering jungle trees, floating mountains, and glowing flora, inhabited by blue-skinned humanoids and diverse wildlife.
The Matrix A simulated city with glass skyscrapers, subway tunnels, and repeating architecture, populated by ordinary humans and hidden agents of the system.
Table 4: We draw on settings inspired by publicly known fictional and real-world contexts spanning both realistic and fantastical domains. Refer to Appendix E for further discussion.
Procedural Content Generation with LLMs. Beyond characters, LLMs have been used for procedural generation of levels, stories, and worlds. For example, Word2World Nasir et al. (2024) generates narratives/world descriptions, Mariogpt Sudhakaran et al. (2023) generates levels, and works in text adventure or interactive story generation Freiknecht and Effelsberg (2020); Hu et al. (2024) show how LLMs can produce dynamic environments. WorldWeaver Jin et al. (2024) also overlaps somewhat, using LLMs to generate world contexts and roles.
Appendix B Dataset Details
Our evaluations draw on two sources of prompts: moral norm statements from Social Chemistry Forbes et al. (2020) and conversational questions from ConvAI2 Dinan et al. (2019). Table 1 summarizes the full set used in our experiments.
Social Chemistry. We use a curated set of everyday moral statements from the Social Chemistry dataset Forbes et al. (2020), which encodes widely held social norms (e.g., “Parents are expected to make sure their kids eat healthy food”). These statements serve as probes for whether generated characters adopt diverse moral stances rather than collapsing into uniform agreement.
ConvAI2. For conversational prompts, we extract candidate utterances from ConvAI2 Dinan et al. (2019) dialogues by filtering for sentences that end with a question mark. From these, we select two categories: general-purpose questions (q1–q5) about hobbies, preferences, or opinions, and sentiment-oriented probes (s1–s5) about feelings or states. This set allows us to examine whether generated characters vary in their interactional responses (e.g., refusal, deflection, compliance) rather than defaulting to always answering.
Appendix C Moral Positions
To model systematic variation in moral reasoning, we define eight moral positions inspired by Moral Foundations Theory Graham et al. (2013). Each position encodes a distinct configuration of stances toward compassion, fairness, loyalty, authority, purity, and liberty.
As summarized in Table 2, the spectrum ranges from strongly prosocial (e.g., prioritizing kindness, justice, and freedom) to highly self-interested (e.g., disregarding others’ suffering, betraying groups, or exploiting authority). Intermediate positions capture more nuanced or context-dependent attitudes, such as conditional cooperation or symbolic rather than binding treatment of purity norms.
Attaching these moral positions to character profiles allows us to probe whether LLM-generated characters exhibit diverse normative judgments rather than collapsing to universally agreeable stances.
Appendix D Interactional Reactions.
In addition to moral stances, PersonaWeaver models conversational behavior through a set of interactional reaction categories (Table 3). These categories capture how characters respond when questioned, ranging from refusals and deflections to volunteering, playful, or hostile replies. By explicitly sampling from this bank, PersonaWeaver breaks the default assistant-like tendency to always comply, instead producing characters that display a broader range of interactional styles reflective of real human dialogue. This variation not only diversifies direct responses but also induces second-order effects in stylistic features, as discussed in Section 2.3.1.
Appendix E Settings
To evaluate character generation at scale, we instantiate populations across a diverse set of narrative contexts inspired by popular TV shows and movies. These settings, listed in Table 4, span both realistic domains (e.g., small towns, urban neighborhoods) and fantastical domains (e.g., magical kingdoms, alien worlds). The split is designed to cover a wide range of cultural, geographic, and stylistic backdrops, ensuring that our analysis of behavioral diversity is not tied to a single environment but instead stress-tested across heterogeneous worlds. Importantly, the prompts do not allude to the original titles but instead describe the physical and social makeup of each world, allowing the model to ground its generations in setting-specific context without relying on surface-level name recognition.
Appendix F Additional Results: LLaMA and Qwen
In the main paper, we presented results using GPT-4o Achiam et al. (2023) as the primary model of study. Specifically, we examined (i) the distribution of moral stances elicited by normative statements (Fig. 3), (ii) the distribution of reactions to open-ended conversational questions (Fig. 4), and (iii) second-order stylistic effects such as filler word usage, punctuation, answer length, and sentiment (Fig. 5). These experiments established the two alignment-induced biases of existing methods and demonstrated how PersonaWeaver mitigates them by inducing broader behavioral and stylistic diversity.
(a) LLaMA: Moral Statements
(b) LLaMA: Reactions to Questions
(c) Qwen: Moral Statements
(d) Qwen: Reactions to Questions
Figure 6: Comparison of Moral and Reaction Biases across two additional models. Top row: LLaMA. Bottom row: Qwen. Each pair contrasts prior work (WorldWeaver Jin et al. (2024), PersonaHub Ge et al. (2024)) with our approach PersonaWeaver. Refer to Section 2.2 and 2.3.1 for discussion.
LLaMA 4 AI (2024)
(a) Filler Words
(b) Punctuations
(c) Answer Length
(d) Sentiment
(e) Filler Words
(f) Punctuations
(g) Answer Length
(h) Sentiment
Qwen 3 Yang et al. (2025)
Figure 7: Comparison of Stylistic Patterns in the generated answers of prior work (WorldWeaver Jin et al. (2024) and PersonaHub Ge et al. (2024)) and our work PersonaWeaver across four stylistic categories (Filler words, Punctuations, Answer Length, and Sentiment). Results are shown for LLaMA 4 AI (2024) (top row) and Qwen 3 Yang et al. (2025) (bottom row). Refer to Section 2.3.1 for further discussion.
Here, we replicate these analyses using two additional frontier models: LLaMA 4 AI (2024) and Qwen 3 Yang et al. (2025). Fig. 6 reports the distributions of moral stances and conversational reactions, while Fig. 7 extends the stylistic comparison across filler words, punctuation, length, and sentiment. Across both models, we observe patterns consistent with those documented for GPT-4o Achiam et al. (2023): prior methods (WorldWeaver Jin et al. (2024), PersonaHub Ge et al. (2024)) collapse into agreement and compliance, while PersonaWeaver broadens the space of moral positions, elicits more varied reactions (e.g., refusals, deflections, playful responses), and induces second-order stylistic diversity.
Taken together, these results suggest that the alignment-induced biases we identify are not model-specific but general across architectures, and that disentangling world-building from behavioral modeling provides a robust means of mitigating them.